OpenStudy (anonymous):

find the solutions to the following system y = x^2 + 3x -4 and y = 4x + 5

OpenStudy (anonymous):

I Dont know how to do this unless i have either y or x

OpenStudy (anonymous):

How about : \(4x + 5 = x^2 + 3x -4 \) does this helps?

OpenStudy (apoorvk):

equate both y's so u have x2 +3x-4 = 4x + 5 simplifies... x2 - x - 9 = 0 solve and ENJOY!!!

OpenStudy (inkyvoyd):

(x+4)(x-5) x=-4, x=5

OpenStudy (across):

The guy is not even here. lol